# Unified Multimodal Interleaved Document Representation for Retrieval

### v2

@@ Appendix A Details of Experimental Setups
Table 9 summarizes the key properties of the datasets used in our experiment, including query modality, target item, entity domain, number of entities, and whether a section ID is provided to indicate the section containing the answer. Additionally, we provide the number of samples in the training, evaluation, and test splits, as well as the size of the corpus. We provide a more detailed explanation of the datasets below.
@@
Appendix B Additional Experimental Results
@@
We analyze the effect of different dataset sizes for training on retriever and reranker performance. To achieve this, we randomly prune samples in the Encyclopedic-VQA dataset at various ratios and report the performance of models trained on these subsets. In Figure 4 (a), we observe that too many samples can degrade retrieval performance. Also, retrieval of textual queries requires fewer samples to reach its optimal performance compared to multimodal retrieval. Similarly, in Figure 4 (b), section retrieval for multimodal queries requires 10% of the dataset to achieve 80% of the full-dataset performance, while section retrieval for textual queries needs only 5%. These observations suggest that additional modalities increase the need for more data. This accounts for the inferior performance of the interleaved format in the ViQuAE experiments (Table 4 (a)). The ViQuAE dataset, at only 2.2% of the size of Encyclopedic-VQA, may be small for the reranker to effectively learn multimodal query-section alignments. We also observe that section retrieval is more challenging, with more samples improving the reranker’s performance. This explains why the ViQuAE reranker has much lower section retrieval scores compared to the one trained on the Encyclopedic-VQA (Table 4 (b)). Given the challenge of obtaining large query-section pair samples, exploring more effective reranker training pipelines is necessary.
Appendix C Case Studies of Document Retrieval
@@
We conduct case studies to demonstrate the advantages of our approach in document retrieval with textual and multimodal queries. In Figure 5 and Figure 6, we illustrate the instances where our approach, which leverages interleaved multimodal contents (e.g., images, tables, and text) within documents, retrieved correct documents for given queries, while the conventional one, which represents documents using only textual data, retrieved documents that appeared to be relevant but were not actually related to the queries.
@@
In Figure 5, a textual query asks for the name of the park located on the north shore of Foster Reservoir. The conventional approach retrieved a document containing unrelated information about a different reservoir. While this document includes terms such as "Peak District National Park" and "North America farm," which make the document superficially relevant, it fails to answer the query. In contrast, our approach identified the document containing the correct answer to the given query.
@@
The advantages of integrating multimodal content into document representation become more apparent in document retrieval with multimodal queries, as shown in Figure 6. For a query consisting of an image of a town hall in Hanover and a textual question about its designer, both our approach and the conventional one retrieved documents about town halls in Germany. However, our approach pinpointed the exact document about the town hall in Hanover, indicating that Hermann Eggert designed the building. The conventional method retrieved a document about a town hall in Munich, which is somewhat related but not an exact match to the query image or question.

### v3

@@ Appendix A Details of Experimental Setups
Table˜8 summarizes the key properties of the datasets used in our experiment, including query modality, target item, entity domain, number of entities, and whether a section ID is provided to indicate the section containing the answer. Additionally, we provide the number of samples in the training, evaluation, and test splits, as well as the size of the corpus. We provide a more detailed explanation of the datasets below.
@@
Appendix B Multi-modality Statistics in Documents
We calculate the statistics related to multi-modality in Wikipedia documents, and find that both images and tables are evenly distributed across the whole documents. To be specific, except for the first section of documents, which contains 1.2 images on average, the distribution of images is consistent across the other sections, containing an average of 0.27 images per section. Also, tables appear less frequently, averaging 0.23 per section, but they are uniformly distributed across all sections.
Appendix C Efficiency of IDentIfy
During the retrieval process, the computational efficiency (i.e., the retrieval latency) of our approach remains the same regardless of the number of interleaved modalities and their compositions, as each document representation (averaged from its section embeddings) is encoded into a fixed-sized vector, whose size is the same as the case where we encode only the text. Also, even if we consider the efficiency within the document embedding process (which is typically not a concern for IR tasks as it can be done offline in parallel), the computational costs and memory usage when embedding multimodal documents are similar to the case of embedding text-only documents, as the factors that impact efficiency are not the number of multimodal content but the number of tokens within documents.
Appendix D Additional Experimental Results
@@
We analyze the effect of different dataset sizes for training on retriever and reranker performance. To achieve this, we randomly prune samples in the Encyclopedic-VQA dataset at various ratios and report the performance of models trained on these subsets. In Figure˜4 (a), we observe that too many samples can degrade retrieval performance. Also, retrieval of textual queries requires fewer samples to reach its optimal performance compared to multimodal retrieval. Similarly, in Figure˜4 (b), section retrieval for multimodal queries requires 10% of the dataset to achieve 80% of the full-dataset performance, while section retrieval for textual queries needs only 5%. These observations suggest that additional modalities increase the need for more data. This accounts for the inferior performance of the interleaved format in the ViQuAE experiments (Table˜4 Right). The ViQuAE dataset, at only 2.2% of the size of Encyclopedic-VQA, may be small for the reranker to effectively learn multimodal query-section alignments. We also observe that section retrieval is more challenging, with more samples improving the reranker’s performance. This explains why the ViQuAE reranker has much lower section retrieval scores compared to the one trained on the Encyclopedic-VQA (Table˜4 Right). Given the challenge of obtaining large query-section pair samples, exploring more effective reranker training pipelines is necessary.
Appendix E Case Study
@@
We conduct case studies to demonstrate the advantages of our approach in document retrieval with textual and multimodal queries. In Figure˜5 and Figure˜6, we illustrate the instances where our approach, which leverages interleaved multimodal contents (e.g., images, tables, and text) within documents, retrieved correct documents for given queries, while the conventional one, which represents documents using only textual data, retrieved documents that appeared to be relevant but were not actually related to the queries.
@@
In Figure˜5, a textual query asks for the name of the park located on the north shore of Foster Reservoir. The conventional approach retrieved a document containing unrelated information about a different reservoir. While this document includes terms such as "Peak District National Park" and "North America farm," which make the document superficially relevant, it fails to answer the query. In contrast, our approach identified the document containing the correct answer to the given query.
@@
The advantages of integrating multimodal content into document representation become more apparent in document retrieval with multimodal queries, as shown in Figure˜6. For a query consisting of an image of a town hall in Hanover and a textual question about its designer, both our approach and the conventional one retrieved documents about town halls in Germany. However, our approach pinpointed the exact document about the town hall in Hanover, indicating that Hermann Eggert designed the building. The conventional method retrieved a document about a town hall in Munich, which is somewhat related but not an exact match to the query image or question.